OpenStudy (anonymous):

log3 x - log3 (2x-1)-3=0

OpenStudy (anonymous):

\[\log_{3} X - \log_{3} (2x-1) -3=0 \]

OpenStudy (anonymous):

Do you want to get the value of x?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

\[\log_{3} x- \log_{3} (2x-1) =3\] \[\log_{3} [x /(2x-1)] =3\log_{3}3\] \[\log_{3} [x /(2x-1)] =\log_{3}3^3\] Since they have the same base, you can 'cancel' the log and get x / (2x-1) = 27 Can you solve it from here?